OpenStudy (anonymous):

Evaluate ∫ (1/2)x^(2) sin(x^3) dx

ganeshie8 (ganeshie8):

Hint : there is a very obvious \(u\) substitution

OpenStudy (anonymous):

-1/6 cos (x^3)

OpenStudy (anonymous):

ok so i got u=x^(3) du=3x^(2) (1/3)du=x^(2) and I got -1/3 cos(x^3)+c . Was -1/3 multiplied nby 1/2 to get -1/6?

OpenStudy (anonymous):

nvm I got it thanks